# Probing Stellar Kinematics with the Time-Asymmetric Hanbury Brown and Twiss Effect

### v1

@@ 1 Introduction
Intensity Interferometry (II) was first pioneered by Hanbury Brown and Twiss (HBT) in the 1950s (Brown & Twiss, 1956), providing measurements of stellar diameters with an accuracy that reaches fractions of a milliarcsecond in the visible (Hanbury Brown et al., 1974). This represented the highest angular resolution achieved in the optical spectrum at the time. After a hiatus of almost half a century, primarily due to limitations in available technology, II has experienced a resurgence in astronomy. One major catalyst for this revival, alongside advances in photon detection technology, was the construction of telescopes with large light-collecting areas and long baselines provided by Imaging Air Cherenkov Telescopes designed for gamma-ray astronomy. Collaborations such as MAGIC (Abe et al., 2024), VERITAS (Acharyya et al., 2024) and H.E.S.S. (Zmija et al., 2024) have successfully measured stellar diameters using II for the first time since the Narrabri Intensity interferometer was used by HBT. But even simpler setups, such as small aperture telescopes equipped with detectors offering high temporal resolution, namely single-photon avalanche photodiodes (SPADs), have enabled the determination of position angles and axial ratios for extended disks surrounding stars (Matthews et al., 2023). In parallel, large-area observational concepts have emerged. Notably, the QUASAR project (Walter et al., 2023) advocates for intensity interferometry systems using telescopes with mirror areas up to $\sim$ 1000 m2 combined with ultra-fast photon counting via SPADs. This approach promises optical angular resolutions at the microarcsecond scale, allowing direct imaging of accretion disks and jets around compact galactic and extragalactic sources.
@@
and the Doppler shift in wavelength is smaller than the optical bandpass
$\Delta\lambda_{\rm kin}<\Delta\lambda$ (5)
On substituting for $\Delta\tau$ and $\Delta\lambda_{\rm kin}$ these two inequalities can be rearranged as
$\frac{v}{c}\;<\;\frac{\Delta\lambda}{\lambda}\;<\;\frac{\lambda}{c\,\Delta t}$ (6)
With recently-developed photon detectors reaching $\Delta t\sim$10\text{\,}\mathrm{ps}$\$ at optical wavelengths, the outer inequality is satisfied for astrophysical sources at $v\sim$10\text{\,}\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$\$. Satisfying the middle inequalities then requires sub-nm bandwidths at optical wavelengths, which is feasible. Exploration of this regime is therefore timely.
@@ 5 Discussion and Conclusion
Our simulations demonstrate that the temporal intensity correlation function is sensitive to the internal kinematics of a light source. We have shown that an ordered, asymmetric LoS velocity distribution, such as that produced by rotation, manifests itself as a time-asymmetric HBT effect. This results in a temporal correlation peak that is skewed or shifted away from a time delay of $\tau=0$. This work establishes a theoretical basis using this asymmetry as a new observable for probing the dynamics of astrophysical systems.
@@
These theoretical predictions also suggest a path toward a laboratory validation. Such experiments would build upon strong foundation of recent work in creating and correlating pseudo-thermal light in tabletop settings to demonstrate the HBT effect, e.g. Dravins et al. (2015) and Nitu Rai et al. (2024). Previous photon-correlation spectroscopy studies have already illustrated the feasibility of detecting narrow-line astrophysical signals in controlled laboratory conditions (Dravins & Germanà, 2008). Its astrophysical applications, from detecting natural lasers to probing coherence properties of light, are reviewed in detail by Dravins (2008). These preceding experiments have largely focused on symmetric temporal correlations arising from broadening caused by thermal motion in scattering fluids. A direct test of our prediction, however, requires demonstrating the asymmetry that arises from an ordered velocity gradient. A more effective experiment design would be to create two pseudo-thermal sources with slight, stable wavelength offset, which serves as a direct and controllable analogue to a kinematic Doppler shift. This could be achieved, for example by passing monochromatic light though a dispersive element, be it prism or diffraction grating to induce a small calibrated wavelength shift. An even more direct method would involve using two tunable lasers, precisely locked to slightly different frequencies.

### v3

@@ 1 Introduction
Intensity Interferometry (II) was first pioneered by Hanbury Brown and Twiss (HBT) in the 1950s (Brown and Twiss, 1956), providing measurements of stellar diameters with an accuracy that reaches fractions of a milliarcsecond in the visible (Hanbury Brown et al., 1974). This represented the highest angular resolution achieved in the optical spectrum at the time. After a hiatus of almost half a century, primarily due to limitations in available technology, II has experienced a resurgence in astronomy. One major catalyst for this revival, alongside advances in photon detection technology, was the construction of telescopes with large light-collecting areas and long baselines provided by Imaging Air Cherenkov Telescopes designed for gamma-ray astronomy. Collaborations such as MAGIC (Abe et al., 2024), VERITAS (Acharyya et al., 2024) and H.E.S.S. (Zmija et al., 2024) have successfully measured stellar diameters using II for the first time since the Narrabri Intensity interferometer was used by HBT. But even simpler setups, such as small aperture telescopes equipped with detectors offering high temporal resolution, namely single-photon avalanche photodiodes (SPADs), have enabled the determination of position angles and axial ratios for extended disks surrounding stars (Matthews et al., 2023). In parallel, large-area observational concepts have emerged. Notably, the QUASAR project (Walter et al., 2023) advocates for intensity interferometry systems using telescopes with mirror areas up to $\sim$ 1000 m2 combined with ultra-fast photon counting via SPADs. This approach promises optical angular resolutions at the microarcsecond scale, allowing direct imaging of accretion disks and jets around compact galactic and extragalactic sources.
@@
while the Doppler shift in wavelength is comparable to the optical bandpass
$\Delta\lambda_{\rm kin}\sim\Delta\lambda$ (5)
On substituting for $\Delta\tau$ and $\Delta\lambda_{\rm kin}$ these two conditions can be rearranged as
$\frac{v}{c}\;\sim\;\frac{\Delta\lambda}{\lambda}\;<\;\frac{\lambda}{c\,\Delta t}$ (6)
With recently-developed photon detectors reaching $\Delta t\sim$10\text{\,}\mathrm{ps}$\$ at optical wavelengths, the outer inequality is satisfied for astrophysical sources at $v\sim$10\text{\,}\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$\$. Satisfying the condition on $\Delta\lambda$ then requires sub-nm bandwidths at optical wavelengths, which is feasible. Exploration of this regime is therefore timely.
@@ 4.3 Results: Kinematic Signatures in Absorption
5 Practical Limitations and Observational Feasibility
The theoretical results presented in this work show that kinematic signatures can be imprinted on the temporal correlation function at the picosecond level. While these results are formally correct, it is crucial to discuss the practical limitations that would occur by an attempt to resolve this. The primary challenges that are up for discussion are atmospheric turbulence and photon statistics.
5.1 The Effect of Atmospheric Turbulence
II is famously insensitive to the first-order phase fluctuations impairing traditional amplitude interferometry. However, the relative arrival times of intensity patterns to observe the HBT effect can be affected by second-order path-length fluctuations. Hanbury Brown (1974) conducted an analysis of the impact that the differential atmospheric path delay between two light paths would have. They concluded that even in extreme cases, the fluctuations were unlikely to exceed 1 ps and would cause negligble loss of correlation.
The original prediction seems to stand in contrast of modern measurements from the field of Lunar Laser Ranging (LLR), which routinely measures timing jitter of about 10 ps (e.g., Ortolani and Barbieri, 2011). This atmospheric noise floor would make it extremely challenging, if not impossible, to resolve the sub-10-picosecond asymmetries predicted in our simulations from the ground. However, it is important to note that the LLR measurement represents a scenario involving a two-way path through the entire atmosphere, with a 2.6 second delay between outgoing and incoming light, during which the atmospheric column changes completely. The differential jitter for an SII measurement, where two nearby telescopes view a star through highly correlated column of air is expected to be smaller, though it remains a fundamental noise source. The $\approx$ 10 picosecond figure should therefore be considered a conservative upper limit for ground-base SII.
Consistent with this expectation, recent measurements of the HBT correlation peak of the Sun (Sliusar et al., 2026) indicate an atmospheric broadening with an upper limit of approximately 6 ps (RMS), supporting the interpretation that the 10 ps LLR value represents a conservative bound for ground-based SII.
The limitation of atmospheric noise could be eliminated altogether by carrying out the observations from space-based instruments like proposed by Klein et al. (2007), where atmospheric effects are absent.
5.2 The Photon Flux Limit
The second major challenge in observing the asymmetric effect pertains to the statistics of photons. We derive an estimate for the two objects for which the model was developed.
For $\gamma$ Cas the apparent magnitude in the R band centered close to the Hα line is 2.3 Wenger et al. (2000). This corresponds to a spectral photon flux density $\Phi$ of roughly $\Phi$ = 7 $\cdot$ 10-6 photons m-2 s-2 Hz-2 in one polarization channel. The emission line has exhibited variability over the past few decades, where recent measurements Neiner et al. (2011) have shown the flux of the spectral line to be 6 times higher in intensity then the the surrounding continuum, therefore $\Phi$ = 4 $\cdot$ 10-5 photons m-2 s-2 Hz-2, same as a star with apparent magnitude 0.5. The Signal-to-Noise Ratio (SNR) in SII can be estimated using
$\mathrm{SNR}=\Phi|V|^{2}A\varepsilon\sqrt{\frac{T}{\Delta t}}$ (9)
where $A$ corresponds to the mirror area of one of the telescopes, $\varepsilon$ gives the optical and detector efficiency and $T$ the observational time. (This formula often appears slightly differently: considering two polarization channels, and replacing $1/\Delta$ with twice the sampling frequency.) Using telescopes with diameter of 1 meter with a total efficiency of 10 $\%$ and observing the spectral line of $\gamma$ Cas for an hour with $\Delta t$ of a picosecond would lead to a SNR of almost 200, with $\Delta t$ = 10 picosecond that would go down to SNR of roughly 60.
In the case of the absorption line, it is to be understood that the intensity would be reduced in comparison with the continuum. This would result in the necessity of longer observation times in order to achieve the same SNR. In particular, (Diamant et al., 2023) measured the spectral flux reduced to a quarter to the relative flux in the continuum going from magnitude 4.0 in the U band to 5.2, leading to SNR of about 2 with the same setup described above with the more optimistic case of $\Delta t$ = 1 ps. This makes the absorption-line measurements significantly more challenging and restricts the feasibility to brighter targets or larger telescopes.
5.3 Parameter Choices and Physical Regimes
Crucially, the $\sim$ 10 ps atmospheric timing jitter is not a fundamental showstopper for the technique. As established in Section 1, the magnitude of the temporal asymmetrx is inversely proportional to velocity. Observation of a slower-moving source will result in a larger and more easily measurable time delay, provided that the spectral filter width is sufficiently small. A source moving at 10 km/s would for example lead to a time delay of orders of tens of picoseconds using a spectral filter with a bandwidth of 0.02 nm.
This shows that the atmospheric limit does not invalidate the technique but rather focuses its application on a specific regime. The higher-velocity regimes and sub-picosecond asymmetries explored in our simulations represent the ideal, atmosphere-free case that could be accessible to a future space-based intensity interferometer, while the larger shifts from lower-velocity sources remain a viable target for ground-based observatories.
@@ 6 Discussion and Conclusion
Our simulations demonstrate that the temporal intensity correlation function is sensitive to the internal kinematics of a light source. We have shown that an ordered, asymmetric LoS velocity distribution, such as that produced by rotation, manifests itself as a time-asymmetric HBT effect. This results in a temporal correlation peak that is skewed or shifted away from a time delay of $\tau=0$ s. This work establishes a theoretical basis using this asymmetry as a new observable for probing the dynamics of astrophysical systems.
@@
These theoretical predictions also suggest a path toward a laboratory validation. Such experiments would build upon strong foundation of recent work in creating and correlating pseudo-thermal light in tabletop settings to demonstrate the HBT effect, e.g. Dravins et al. (2015) and Nitu Rai et al. (2024). Previous photon-correlation spectroscopy studies have already illustrated the feasibility of detecting narrow-line astrophysical signals in controlled laboratory conditions (Dravins and Germanà, 2008). Its astrophysical applications, from detecting natural lasers to probing coherence properties of light, are reviewed in detail by Dravins (2008). These preceding experiments have largely focused on symmetric temporal correlations arising from broadening caused by thermal motion in scattering fluids. A direct test of our prediction, however, requires demonstrating the asymmetry that arises from an ordered velocity gradient. A more effective experiment design would be to create two pseudo-thermal sources with slight, stable wavelength offset, which serves as a direct and controllable analogue to a kinematic Doppler shift. This could be achieved, for example by passing monochromatic light though a dispersive element, be it prism or diffraction grating to induce a small calibrated wavelength shift. An even more direct method would involve using two tunable lasers, precisely locked to slightly different frequencies.
@@ Acknowledgments
Additionally we would like to thank the anonymous Referee for their helpful comments and feedback, which greatly strengthened the overall manuscript.
This research has made use of the SIMBAD database, CDS, Strasbourg Astronomical Observatory, France and of the BeSS database, operated at LESIA, Observatoire de Meudon, France: http://basebe.obspm.fr